OpenStudy (anonymous):

Find the point(s) on the parabola y=x^2 which is/are closest to the point (0,2). Fully justify your answer using calculus.

OpenStudy (dumbcow):

you want to minimize distance from parabola to point using distance formula \[d = \sqrt{(x-0)^2 + (y-2)^2} = \sqrt{x^2 + (x^2 -2)^2}\] set derivative equal to 0 solve for x

OpenStudy (anonymous):

I got the points (-sqrt{3/2},3/2) and (sqrt{3/2},3/2) Would these be correct?

OpenStudy (dumbcow):

yes thats correct

OpenStudy (anonymous):

Yay thank you!